Given x + iy form, write in polar form: 5-i
Answer by fractalier(2101)

The r is found by sqrt(x^2 + y^2) and the angle is the arctan(y/x)...so we get
5 - i = (sqrt(26), arctan(-1/5)) = (5.1, -11.3°)